# Quantitative evaluation of methods to analyze motion changes in single-particle experiments

### v2

@@ Appendix A Simulations of diffusion and interaction models
Simulations are performed considering particles diffusing in a square box of size $L$ with reflecting boundary conditions. However, to avoid boundary effects, the fields of view used for the challenge datasets correspond to a square region of size $L_{\rm FOV}\ll L$ within the central part of the original box (2(b)).
For Track 1, trajectory coordinates are used as sub-pixel localizations of individual particles to simulate movie frames as in single-molecule fluorescence experiments [5]. Each particle has a random intensity $I_{i}$ that corresponds to the total number of photons collected by the detector. $I_{i}$ is drawn from a uniform distribution in the interval $[I_{\rm min},I_{\rm max}]$ and fluctuates over time according to a normal distribution with mean $I_{i}$ and standard deviation $\sigma_{I}$. Each particle is rendered as a diffraction-limited spot using an Airy disk as a point-spread function (PSF) with full width at half maximum ${\rm FWHM_{\rm PSF}}=2.1$ px. A constant background of $I_{\rm bg}=100$ counts is added to each frame. Images are corrupted with Poisson noise.
For Track 2, trajectory coordinates are corrupted with noise from a Gaussian distribution with zero mean and standard deviation $\sigma_{N}$ to take into account the finite localization precision obtained in tracking experiments.
@@
  • $[K_{1},K_{2},\dots,K_{n}]$: average values of the (Gaussian) distribution of the generalized diffusion coefficient for each of the $n$ diffusive states considered in a given experiment, with support $[10^{-12},10^{6}]$ ${\rm pixel^{2}\cdot frame^{-\alpha}}$.
  • $[\sigma_{K_{1}},\sigma_{K_{2}},\dots,\sigma_{K_{n}}]$: standard deviations of the (Gaussian) distribution of the generalized diffusion coefficient for each of the $n$ diffusive states considered in a given experiment. If not provided, it is considered to be equal to 0 (i.e., the distribution is $\delta(K-K_{i})$).
  • $[\alpha_{1},\alpha_{2},\dots,\alpha_{n}]$: average values of the (Gaussian) distribution of the anomalous diffusion exponent for each of the $n$ diffusive states considered in a given experiment, with support $(0,2)$.
  • $[\sigma_{\alpha_{1}},\sigma_{\alpha_{2}},\dots,\sigma_{\alpha_{n}}]$: standard deviations of the (Gaussian) distribution of the anomalous diffusion exponent for each the $n$ diffusive states considered in a given experiment. If not provided, it is considered to be equal to 0 (i.e., the distribution is $\delta(\alpha-\alpha_{i})$).
  • $L$: size of the box where trajectories are simulated with reflecting boundary conditions.
@@
  • $T$: duration of the recording over each FOV, given as the number of time steps $\Delta t$. It also corresponds to the maximum trajectory duration. For the challenge, we set $T=500$;
@@
    $SNR=\frac{I_{\rm peak}}{\sqrt{I_{\rm peak}+I_{\rm bg}}}.$ (14)
@@
  • $t_{\rm min}$: minimum distance between changepoints, corresponding to the minimum amount of time that a particle spends in a state. Shorter segments are eliminated by smoothing the time trace of the state label using a majority filter with a window of 5 steps. For the challenge, we set $t_{\rm min}=3$.;
A schematic representation of each of the models presented below is shown in 2(a).
A.1 Model 1 - Single-state model (SSM)
@@
A.2 Model 2 - Multi-state model (MSM)
The multi-state model is a Markov model describing particles undergoing FBM whose diffusion properties can change at random times. The number of states $S$ is fixed for a given experiment as are the parameters defining the distributions of $K$ and $\alpha$ for each state. For each trajectory, $S$ values of $\alpha$ and $S$ values of $K$ are sampled from the distribution of the corresponding states, i.e., one per state. At every time step, a diffusing particle has a given probability to undergo a change in one of its diffusive parameters (either $\alpha$ or $K$). The probability of switching is given by a transition matrix $M$. Namely, $M_{i,j}$ is the probability of switching from state $i$ to state $j$ at each time step. In the same sense, $M_{i,i}$ is the probability of remaining in state $i$. The residence time in a given state $i$ can be directly calculated from the previous probability as
@@ A.2 Model 2 - Multi-state model (MSM)
[…]
A.3 Model 3 - Dimerization (DIM)
@@
[…]
A.4 Model 4 - Transient-confinement model (TCM)
This model considers an environment with $N_{\rm c}$ circular compartments of radius $r_{\rm c}$. The compartments are distributed randomly throughout the environment such that they do not overlap. We consider that the compartments are osmotic, i.e., a particle reaching their boundary from the exterior has a probability 1 of entering them, but a particle reaching the boundary from the interior of a compartment has a probability $T$ of exiting it (and $1-T$ of being reflected back to the interior of the compartment). The diffusion inside and outside the compartment is different, hence defining two diffusive states. For each trajectory, two values of $\alpha$ and two values of $K$ are sampled from the corresponding distributions, corresponding to the motion outside and inside the compartments.
[…]
A.5 Model 5 - Quenched-trap model (QTM)
@@
[…]
List of acronyms
RMSE
root mean square error
MAE
mean absolute error
MSLE
mean squared logarithmic error
JSC
Jaccard similarity coefficient
FP
false positive
TP
true positive
FN
false negative
CP
changepoint
FOV
Field of view
SSM
Single-state model
MSM
Multi-state model
DIM
Dimerization model
TCM
Transient-confinement model
QTM
Quenched-trap model

### v3

@@ 6.1 Simulations of diffusion and interaction models
Simulations are performed considering particles diffusing in a square box of size $L$ with reflecting boundary conditions. However, to avoid boundary effects, the fields of view used for the Challenge datasets correspond to a square region of size $L_{\rm FOV}\ll L$ within the central part of the original box (Fig.˜2b).
For Track 1, trajectory coordinates are used as sub-pixel localizations of individual particles to simulate movie frames as in single-molecule fluorescence experiments [5]. Each particle has a random intensity $I_{i}$ that corresponds to the total number of photons collected by the detector. $I_{i}$ is drawn from a uniform distribution in the interval $I_{\rm min},I_{\rm max}]$ and fluctuates over time according to a normal distribution with mean $I_{i}$ and standard deviation $\sigma_{I}$. Each particle is rendered as a diffraction-limited spot using an Airy disk as a point-spread function (PSF) with full width at half maximum ${\rm FWHM_{\rm PSF}}=2.1$ px. A constant background of $I_{\rm bg}=100$ counts is added to each frame. Images are corrupted with Poisson noise.
For Track 2, trajectory coordinates are corrupted with noise from a Gaussian distribution with zero mean and standard deviation $\sigma_{N}$ to take into account the finite localization precision obtained in tracking experiments. All simulated trajectories were generated without missing frames: no gaps were introduced, yielding continuous tracks to isolate segmentation performance from linking or gap-filling complexities.
@@
  • $K_{1},K_{2},\dots,K_{n}]$: average values of the (Gaussian) distribution of the generalized diffusion coefficient for each of the $n$ diffusive states considered in a given experiment, with support $10^{-12},10^{6}]$ $\rm pixel^{2}/frame^{\alpha}$}.
  • $\sigma_{K_{1}},\sigma_{K_{2}},\dots,\sigma_{K_{n}}]$: standard deviations of the (Gaussian) distribution of the generalized diffusion coefficient for each of the $n$ diffusive states considered in a given experiment. If not provided, the standard deviation is considered to be equal to 0 (i.e., the distribution is $\delta(K-K_{i})$).
  • $\alpha_{1},\alpha_{2},\dots,\alpha_{n}]$: average values of the (Gaussian) distribution of the anomalous diffusion exponent for each of the $n$ diffusive states considered in a given experiment, with support $(0,2)$.
  • $\sigma_{\alpha_{1}},\sigma_{\alpha_{2}},\dots,\sigma_{\alpha_{n}}]$: standard deviations of the (Gaussian) distribution of the anomalous diffusion exponent for each of the $n$ diffusive states considered in a given experiment. If not provided, the standard deviation is considered to be equal to 0 (i.e., the distribution is $\delta(\alpha-\alpha_{i})$).
  • $L$: size of the box in which trajectories are simulated with reflecting boundary conditions.
@@
  • $T$: duration of the recording over each FOV, given as the number of time steps $\Delta t$. It also corresponds to the maximum trajectory duration. For the Challenge, we set $T=200$;
@@
    $\mathrm{SNR}=\frac{I_{\rm peak}}{\sqrt{I_{\rm peak}+I_{\rm bg}}}.$ (15)
@@
  • $t_{\rm min}$: minimum distance between changepoints, corresponding to the minimum amount of time that a particle spends in a state. Shorter segments are eliminated by smoothing the time trace of the state label using a majority filter with a window of 5 steps. For the Challenge, we set $t_{\min}=3$ frames to test the sensitivity and robustness of the segmentation methods under minimal data conditions.
A schematic representation of each of the models presented below is shown in Fig.˜2a.
6.1.1 Model 1 - Single-state model (SSM)
@@
6.1.2 Model 2 - Multi-state model (MSM)
The multi-state model is a Markov model describing particles undergoing FBM whose diffusion properties can change at random times. The number of states $S$ is fixed for a given experiment as are the parameters defining the distributions of $K$ and $\alpha$ for each state. For each trajectory, $S$ values of $\alpha$ and $S$ values of $K$ are sampled from the distribution of the corresponding states, i.e., one per state. At every time step, a diffusing particle has a given probability to undergo a change in one of its diffusive parameters (either $\alpha$ or $K$). The probability of switching is given by a transition matrix $M$. Namely, $M_{ij}$ is the probability of switching from state $i$ to state $j$ at each time step. In the same sense, $M_{ii}$ is the probability of remaining in state $i$. The residence time in a given state $i$ can be directly calculated from the previous probability as
@@ 6.1.2 Model 2 - Multi-state model (MSM)
Model 2 (MSM) parameters
  • $M$: transition matrix between diffusive states.
6.1.3 Model 3 - Dimerization (DIM)
@@
Model 3 (DIM) parameters
  • NN: number of diffusing particles in the box of size $L$.
  • $r$: interaction radius, corresponding to the radius of the diffusing particles.
  • $P_{\rm b}$: probability that two particles bind to form a dimer in each time step. For this to happen, the particles must be at a distance $d<2r$.
  • $P_{\rm u}$: probability that a dimer breaks up at each time step so that the two particles go back to diffusing independently.
6.1.4 Model 4 - Transient-confinement model (TCM)
This model considers an environment with $N_{\rm c}$ circular compartments of radius $r_{\rm c}$. The compartments are distributed randomly throughout the environment such that they do not overlap. We consider that the compartments are osmotic, i.e., a particle reaching their boundary from the exterior has a probability 1 of entering them, but a particle reaching the boundary from the interior of a compartment has a probability $T$ of exiting it (and 1T1-T of being reflected back to the interior of the compartment). The diffusion inside and outside the compartment is different, hence defining two diffusive states. For each trajectory, two values of $\alpha$ and two values of $K$ are sampled from the corresponding distributions, representing the motion outside and inside the compartments.
Model 4 (TCM) parameters
  • $N_{\rm c}$: number of compartments in the box of size $L$.
  • $r_{\rm c}$: radius of the compartments.
  • $T$: transmittance of the boundary. Probability that a particle reaching the boundary from inside the compartment exits the compartment.
6.1.5 Model 5 - Quenched-trap model (QTM)
@@
Model 5 (QTM) parameters
  • $N_{\rm t}$: number of traps in the box of size $L$.
  • $r_{\rm t}$: radius of the traps.
  • $P_{\rm b}$: probability that a particle binds to a trap and gets immobilized. For that to happen, a particle must be at a distance $d<r_{\rm t}$ from the trap.
  • $P_{\rm u}$: probability that a trapped particle unbinds from a trap and starts diffusing independently at each time $\delta t$.
6.2 Dataset structure
The datasets used in the Challenge (Supp. Fig. 9) include different experiments, each contained in a folder labeled with a sequential number (EXP_[exp number]) and corresponding to a specific model and a fixed set of parameters. The information about the model and the parameters is unknown to Challenge participants. Each experiment folder contains a list of files labeled with a sequential number (FOV_[fov number]) associated with 30 FOVs. Each FOV reports data from a variable number of particles diffusing on a $128\times 128$ pixel2 area.
For the Video Track, the coordinates of the particles in the same FOV are used to generate 200-frame videos as a series of 8-bit images in the multi-tiff format using Deeptrack 2.1 [5]. Noise is added to the synthetic images to account for background fluorescence and shot noise. A map corresponding to the segmentation of VIP particles at the first frame for which CPs and diffusion parameters must be detected is also provided as a tiff file. Connected components of the map are labeled with unique integer values that correspond to the particle index.
For the Trajectory Track, we provide a csv file for each FOV with a table whose columns contain trajectory index, time step, $x$-coordinate, and $y$-coordinate. Coordinates of simulated trajectories are corrupted with Gaussian noise corresponding to finite (subpixel) localization precision. The trajectories have a maximum length of 200 frames.
Exemplary data for all the models are shown in Supp. Fig. 10. Files in different Tracks labeled with the same experiment and the FOV index (e.g., Track_1/EXP_4/FOV_3.tiff and Track_2/EXP_4/FOV_3.csv) include simulations obtained with the same set of dynamics parameters but do not correspond to the motion of the same set of particles.
7 Data availability
The labeled benchmark dataset used in this study is available on Zenodo [94]. All datasets generated for the Challenge can be accessed on the Codalab platform (registration required). Source data for all figures are provided with this paper.
8 Code availability
All code used to generate the Challenge datasets is publicly available via the andi_datasets repository on GitHub: https://github.com/AnDiChallenge/andi_datasets [71].